OpenStudy (anonymous):

3x+16y=6 what is x and y

OpenStudy (kira_yamato):

I need another equation... The relationship between x and y is x = 2 - 16y/3

OpenStudy (anonymous):

\[15x+40y=32\]

OpenStudy (anonymous):

Use the value for X that Kira got and plug it into the second equation, then solve. 15(2 - 16y/3) + 40y = 32 30 - 80y + 40y = 32 -40y = 2 y = -1/20 Then plug the value for Y into the first equation. x = 2 - 16(-1/20)/3 x = 4 2/5

OpenStudy (anonymous):

That's 4 and 2/5 or 4.4

OpenStudy (kira_yamato):